Prove that any n-th root of unity is a primitive d-th for a uniqued/n ?


1
Expert's answer
2021-06-07T10:33:22-0400

Let F

F be a field and let n∈N

 and let ε∈F

 (F is an algebraic clausure of F) be a 

nth root of unity (

ε is a root of X

n−1∈F[X])

"X_n\u22121\u2208F[X]" . Can I state that there exists d∈N

d∈N, with d|n, such that ε is a d

dth primitive root of unity? (

ε is d-th primitive root of unity if ε

ε is d-th root of unity and ε generates all the elements in the group of d

d-th roots of unity, which is a subgroup in "F\n\n_{0}" . I am going to denote by

Wk the group of kth roots of unity and W∗k the elements in 

Wk which are primitive.

I have not found any counter example. For instance, if we take 

F=C, then 

W4={1,i,−1,i}. In this group, the only elements which do not belong to W∗3

 are 1

1 and −1

−1; however, −

−1∈W4∗ and 1∈W


1∈W1∗, so no counterexample can be found here. On the other hand, if we take 

F as the finite field of 

9 elements, 

F=GF(9) (whose operations between the elements can be found


here: http://www.cs.miami.edu/home/burt/learning/Csc609.032/notes/gf9example.html),


we can see that

W4={1,2,x,2x}. Here, 

W4∗={x,2x}, but 

2∈W2∗ and 

1∈W1∗. Finally and it is my last example, if we take

F=GF(4), we have that 

X4−1=(X−1)4 and 

X2−1=(X−1)2, so 1

1 is a primitive element, as 

W4=W2={1}.

My attempt for the proof: let d

d the order for ε

ε as element in Wn

Wn (each element in a group generates a cyclic group), which is a cyclic and finite group. Then 

εd=1 and ⟨ε⟩={1,ε,…,εd−1}. Apparently, we are done because 

εd=1, which shows that 

ε∈Wd, but (εj)d=1j=1 for all j=0,1,…,d−1

j=0,1,…,d−1, which shows that 

⟨ε⟩⊂Wd and 

|Wd|≤d, which shows that 

⟨ε⟩=Wd.

The main problem which I have found in the proof is that I cannot state that the group 

Wn has n elements (see my last example), it happens in finite groups or in groups which characteristic is not zero. Generally, I can state that 

m≤d (it is an equality in C

C por example), where 

m=|Wd| but I cannot state that m=d

m=d.


hence proved.